# A Geometric Theory of Cognition

### v1

@@ Abstract
A central challenge in understanding cognition is that prevailing theories impose strong assumptions on the nature of internal mental representations. Probabilistic accounts characterise cognition as approximate inference; [11] neural theories emphasise dynamics in recurrent circuits; [12, 13] symbolic approaches posit discrete, rule-based structures; and contemporary machine-learning models rely on particular function classes, optimisation objectives or architectural constraints. Although these approaches have achieved notable successes, their representational commitments make it difficult to explain how diverse cognitive phenomena—such as intuitive, habitual responses and slower, deliberative reasoning—emerge from shared underlying principles. Dual-process theories describe the coexistence of fast and slow modes of thought, [14] but do not account for why such regimes arise or how they are related within a single computational system.
@@ Results
First, the framework unifies a wide range of existing cognitive and computational theories. Bayesian inference, predictive coding, reinforcement learning, deep representation learning, the free-energy principle, and dual-process accounts of intuition and deliberation all emerge as special cases of gradient flow under different choices of potential $J$ and metric $G$. This provides a mathematical explanation for the empirical success of these models and clarifies their relationships: rather than being competing theories, they can be interpreted as different parameterisations of a single underlying geometric principle governing learning and decision dynamics.
@@
Third, the framework offers a principled basis for designing artificial general intelligence. Representing an agent’s internal configuration as a latent cognitive state $\eta$, updated through the gradient flow of a unified potential $J$, subsumes both deep learning [25, 26] and reinforcement learning [27, 28, 29] within a common geometric description. Learning corresponds to shaping the potential and metric through data, while decision-making corresponds to following the induced flow. This perspective avoids modular architectures separating memory, prediction, reward, and planning. Many recent approaches—world models, [30, 31] JEPA-style self-supervised systems, [32] GFlowNets, [33] and differentiable decision-making [34] pipelines—can be viewed as concrete instances of this formulation. The anisotropic metric naturally induces fast intuitive and slow deliberative processes, enabling agents to allocate computational resources adaptively and to reason over multiple time scales.

### v2

@@ Abstract
A central challenge in understanding cognition is that prevailing theories impose strong assumptions on the nature of internal mental representations. Probabilistic accounts characterise cognition as approximate inference; [11] neural theories emphasise dynamics in recurrent circuits; [12, 13] symbolic approaches posit discrete, [14] rule-based structures; [15] and contemporary machine-learning models rely on particular function classes, optimisation objectives or architectural constraints. [16] Although these approaches have achieved notable successes, their representational commitments make it difficult to explain how diverse cognitive phenomena—such as intuitive, habitual responses and slower, deliberative reasoning—emerge from shared underlying principles. Dual-process theories describe the coexistence of fast and slow modes of thought, [17] but do not account for why such regimes arise or how they are related within a single computational system.
@@ Results
First, the framework unifies a wide range of existing cognitive and computational theories. Bayesian inference, [28]predictive coding, [29] reinforcement learning, [30] deep representation learning, [31] the free-energy principle, [32] and dual-process accounts of intuition and deliberation [33] all emerge as special cases of gradient flow under different choices of potential $J$ and metric $G$. This provides a mathematical explanation for the empirical success of these models and clarifies their relationships: rather than being competing theories, they can be interpreted as different parameterisations of a single underlying geometric principle governing learning and decision dynamics.
@@
Third, the framework offers a principled basis for designing artificial general intelligence. Representing an agent’s internal configuration as a latent cognitive state $\eta$, updated through the gradient flow of a unified potential $J$, subsumes both deep learning [34] and reinforcement learning [35, 36] within a common geometric description. Learning corresponds to shaping the potential and metric through data, while decision-making corresponds to following the induced flow. This perspective avoids modular architectures separating memory, prediction, reward, and planning. Many recent approaches—world models, [37, 38] JEPA-style self-supervised systems, [39] GFlowNets, [40] and differentiable decision-making [41] pipelines—can be viewed as concrete instances of this formulation. The anisotropic metric naturally induces fast intuitive and slow deliberative processes, enabling agents to allocate computational resources adaptively and to reason over multiple time scales.